### v1

# FinPos: A Position-Aware Trading Agent System for Real Financial Markets

Abstract.
The exceptional potential of large language models (LLMs) in handling text information has garnered significant attention in the field of financial trading. However, current trading agents primarily focus on single-step trading tasks and lack awareness of continuous position management. Therefore, we propose a position-aware trading task designed to simulate a more realistic market. To address this task, we develop a trading agent system, FinPos, optimized for position management. FinPos is able to interpret various types of market information from a professional perspective, providing a reliable basis for positioning decisions. To mitigate the substantial market risks arising from position fluctuations, FinPos employs dual decision agents. Furthermore, the continuous nature of position management necessitates our adoption of multi-timescale rewards, which in turn empowers FinPos to effectively balance short-term fluctuations against long-term trends. Extensive experiments demonstrate that FinPos surpasses state-of-the-art trading agents in the position-aware trading task, which closely mirrors real market conditions. More importantly, our findings reveal that LLM-centered agent systems exhibit a vast, largely unexplored potential in long-term market decision-making.
Key words and phrases:
Trading Agent, Position Aware, Real Markets.
Corresponding author.
1. Introduction
Figure 1. With the introduction of position awareness, the agent must not only predict current market trends but also manage the remaining positions in the account. Agents developed for tasks without position awareness are inadequate for addressing the new challenges posed by position-aware trading tasks.
As communication and information technologies evolve, the financial markets are seeing an unprecedented surge in critical, market-moving information Revsine et al. (2021). Due to inherent cognitive limitations, human traders struggle to process the sheer volume of financial data efficiently Black (1986), which poses a serious challenge for real-time market responses. As a result, automated trading systems Huang et al. (2019); Tran et al. (2023); Adediran et al. (2025), which leverage sophisticated models, are becoming increasingly vital within the financial industry.
In the past decade, deep reinforcement learning (DRL) agents Fischer (2018); Shakya et al. (2023) gained considerable attention due to their ability to handle large financial datasets. These agents utilize deep neural networks Millea (2021) to extract meaningful features, enabling them to approximate complex market dynamics. While DRL agents can learn strategies to maximize investment returns, they still struggle to incorporate unstructured textual information—such as news articles Lillo et al. (2015) and financial reports Oyewole et al. (2024). Furthermore, their limited interpretability Balhara et al. (2022) makes it difficult for human traders to trust or collaborate with them effectively.
Recently, large language models (LLMs) Hurst et al. (2024); Liu et al. (2024); Yang et al. (2025) have demonstrated remarkable potential in handling text-modal information and solving complex tasks, such as reasoning and decision-making Zhai et al. (2025). This suggests that LLM-centered agent systems have the potential to transcend the inherent capability boundaries of traditional DRL trading models. Specialized LLM agents designed for financial trading, like FinGPT Yang et al. (2023), FINMEM Yu et al. (2024a), and FinAgent Zhang et al. (2024b), have already achieved groundbreaking results. However, as shown in the left panel of Fig. 1, current LLM agents primarily tackle single-step trading tasks, making them more akin to isolated simulations than realistic financial decision-making systems. In real markets, balancing risk and return requires a sophisticated trading agent to shift its focus from short-term market fluctuations to long-term position management. Therefore, we propose a position-aware trading task, which more closely aligns with real-world trading markets than previous setups.
Existing agent architectures, as shown in Fig. 1, struggle to sustain high yield rates when tasked with position management. This is mainly because they lack three essential capabilities: (1) Position-based risk management: When an agent is responsible for determining investment scale, the associated risks of its decisions are considerably greater than those encountered in daily trend forecasting tasks. (2) Long-term planning capability: Trading tasks that are re-initialized daily are prone to developing agents with a myopic trading perspective. Effective position management, however, necessitates a more long-term investment strategy. (3) In-depth market analysis ability: Short-term trends are often identifiable through market sentiment, yet a much deeper analysis is essential for understanding long-term growth potential. To address these challenges, we propose FinPos, a trading agent capable of managing portfolio positions within real financial markets.
To address position-based risk management, FinPos adopts a multi-agent decision-making system that separates the direction decision agent (buy/sell/hold) from the scale decision agent (transaction volume). This decoupling makes risk management more flexible and controllable. The investment direction decision agent is exclusively tasked with determining whether to buy, sell, or hold on each trading day. Meanwhile, the investment scale decision agent dictates the specific transaction volume to buy or sell, subject to an upper limit on single transaction sizes to ensure capital safety. Additionally, we incorporate risk-awareness and position-management awareness into the investment scale decision agent via prompts, keeping investment risk within controllable bounds. More importantly, as financial market risk originates not only from firm-specific events but also from macroeconomic developments (e.g., tariff regimes), FinPos incorporates macroeconomic news monitoring to proactively hedge against systemic risks.
Long-term planning capability is an essential requirement for a position-aware trading agent. Current financial LLM agents are limited by daily trading tasks, leading them to exclusively focus on the next day’s price movements. This limitation results in agents adopting narrow, day-to-day trading strategies. In real markets, agents requires not only the ability to anticipate short-term fluctuations but also a long-term perspective on assets. To address this, we design a comprehensive short-term and long-term reward framework for FinPos. First, we employ a window-based design that evaluates today’s decisions against outcomes over multiple days, ensuring that short-term actions are judged in a longer context. Second, we embed a multi-timescale reward into the reflection process, so that after each action the agent re-examines its choice and measures consistency with both immediate and cumulative returns. Together, these components allow FinPos to move beyond “one-day-at-a-time” strategies and acquire genuine long-term planning ability during training.
By observing the reasoning chain of trading agents, we observe that current trading agents often analyze market information superficially and lack causal reasoning abilities. For instance, when trading Tesla stock, the agent may overestimate the relevance of news such as ”Musk invests in company A” while dismissing significant news like ”major changes in U.S. tariff policy” as irrelevant to trading decisions. This issue arises because the information analysis agent lacks fundamental common sense about the specific forms of information. To enhance in-depth market analysis ability, we employ a strategy of independent analysis by multiple agents for various types of market information (e.g., financial reports, news) and inject fundamental domain-specific knowledge and focused guidance on key issues into each analyst. With this mechanism, FinPos achieves deeper insights into the market.
We conduct extensive experiments using data from multiple real stocks to demonstrate the effectiveness of the FinPos agent in authentic market environments. FinPos outperforms several state-of-the-art (SOTA) financial agents and is the first agent equipped with investment position management capabilities. Moreover, a series of ablation studies underscored the critical importance of risk management, long-term planning, and in-depth market analysis for the agent’s ability to manage investment positions effectively.
2. Related Works
2.1. LLM-Based AI Agents
As researchers increasingly recognize the extraordinary capabilities of large language models (LLMs) in tasks such as summarization Zhang et al. (2024a), reasoning Ke et al. (2025), and decision-making Eigner and Händler (2024), AI agents built around LLMs are gaining significant attention. These LLM agents operate as closed-loop systems, autonomously gathering information, storing memories, reasoning, making decisions, and observing environmental feedback. The advancement of LLMs is propelling AI agents into diverse fields, from open-ended artistic creation to precise medical applications. For instance, MusicAgent Yu et al. (2023) combines various music-related tools and autonomous workflows to assist artists in music creation. MedAgents Wang et al. (2025) use a multi-disciplinary collaboration framework to improve the efficiency of medical diagnosis. Programmer Agent automates the development of software projects. ResearchAgent Li et al. (2024b) helps researchers generate innovative ideas from extensive literature reviews. Personal assistant agents Li et al. (2024a) engage closely with users, providing solutions for tasks and emotional support. These applications illustrate the vast potential of LLM agents in multiple facets of human society. We aim to bring this potential to the financial markets, making LLM agents key players in investment activities.
2.2. Financial LLM-Based Agents
The evolution of financial agent architectures has exhibited significant progress through successive innovations and expansions in cognitive modeling and system integration. FINMEM Yu et al. (2024a) was the first to establish a fundamental trading agent architecture centered around analysis, memory, and decision modules. Building upon this foundation, FinAgent Zhang et al. (2024b) and FinVision Fatemi and Hu (2024) integrate multimodal data—encompassing numerical, textual, and visual information—to analyze both dynamic financial market trends and historical trading patterns. FINCON Yu et al. (2024b) and TradingAgents Xiao et al. (2024) introduced the concept of investment group architectures into agent systems, employing multi-agent collaboration mechanisms to enhance decision-making reliability. Fundamentally, these agents primarily address a simplified financial game focused on predicting next-day price movement. In real markets, however, agents must not only needs to determine the investment direction, but also decide when to adjust exposure and how much to allocate, considering a comprehensive historical investment context. Therefore, this paper adopts a position-aware trading task that more closely approximates real market environments, enabling a more authentic validation of trading agents’ capabilities.
3. Task Definition
3.1. Single-Step Trading Task
We define a baseline environment as a Single-Step Trading Task, where the agent’s position is automatically liquidated at each timestep, and no position is carried over. Within this framework, the agent makes a trading decision $a_{t}\in\{-1,0,1\}$ at each timestep $t$, corresponding to selling, holding, or buying, respectively. Each decision assumes immediate liquidation at the next timestep, thereby making each action independent, with no continuation of position states. The cumulative return $R$ is the sum of the logarithmic returns at each step:
$R=\sum_{t=1}^{N}r_{t},\quad r_{t}=a_{t}\times\log\frac{\text{price}[t+1]}{\text{price}[t]}$ (1)
This setup is widely adopted in studies such as FINMEM Yu et al. (2024a) due to its simplification of trading processes, which stabilizes the training environment and allows more straightforward model evaluation. However, this simplification comes with significant limitations: it completely ignores the continuity of holding positions and the gains and losses arising from holding positions in real trading, thus cannot capture the associated compound returns or the risk exposure carried across periods. Without modeling positions, the agent has no basis for risk control or for detecting market trends, both of which are central to real trading.These omissions weaken the agent’s learning and generalization in essential skills such as position management and risk control.More critically, by reducing trading to short-term, step-by-step predictions, such a setup underutilizes the natural strengths of LLMs in reasoning over longer horizons. Consequently, advancing financial trading agents towards practical application requires a task that more closely mirrors real market conditions.
3.2. Position-Aware Trading Task
Compared to the aforementioned simplified setup, this paper proposes a trading task configuration that aligns more closely with real market behavior, Position-Aware Trading Task. In this environment, the agent makes a trading decision on each trading day (each timestep $t$ based on the closing price) and explicitly maintains its position state. The agent’s decisions no longer involve automatic liquidation; rather, it must consider the continuation and adjustment of current positions in future timesteps. Returns $R$ are calculated from the cumulative logarithmic returns based on position:
$R=\sum_{t=1}^{N}r_{t},\quad r_{t}=\text{position}_{t}\times\log\frac{\text{price}[t+1]}{\text{price}[t]}$ (2)
Where $\text{position}_{t}$ represents the agent’s current position state, dynamically updated based on past trading decisions. This setup not only more accurately models the continuity and trend dynamics of holding positions in trading but also allows the agent to naturally incorporate key capabilities in strategy learning, such as position control, risk exposure management, and entry-exit timing. The Position-Aware Trading Task provides the agent with a more challenging and practically meaningful learning objective, offering stronger potential for strategy generalization and real-world applicability.
Additionally, this task better aligns with the capability characteristics of trading LLM agents. Compared to traditional quantitative models, LLMs excel at extracting long-term trends and underlying causal structures from complex semantic information rather than performing high-frequency precise numerical optimization. This capability gives them greater potential in mid-to-long-term strategy formulation. Under the Single-Step Trading Task setup, the agent primarily focuses on the price direction of the next step, which somewhat forces the LLM struggling to build coherent strategies or cognition. By introducing position continuity and strategy coherence, the Position-Aware Trading Task provides LLM agents with a more fitting testing platform, fully unlocking their potential in long-term decision-making and complex information integration. Furthermore, this setup enables agents to construct an internally consistent position management logic centered around fundamentals, market sentiment, and macro semantic information. The combination of long-term and semantic-driven strategies is precisely where LLMs hold their advantages in financial trading tasks.
4. Architecture of FinPos
The architecture of FinPos (Fig. 2) consists of three core modules: the market signal processing and analysis module, the trading decision module, and the multi-timescale reward reflection module. The signal processing module introduces a two-level agent hierarchy—filtering and analysis—to process raw data, with the results stored in a memory layer. The decision module advances beyond prior research that only predicts direction by introducing a position-aware mechanism. A direction decision agent determines the trading orientation, while a quantity decision agent dynamically adjusts position size in response to market risk exposure. The reward module adopts a multi-timescale reward feedback mechanism to enhance the long-term behavioral stability of the model and its responsiveness to trading opportunities. The details of these three modules are presented in the subsequent sections, while all agent prompts are provided in the Appendix A.
Figure 2. Architectural Details of FinPos: Initially, multiple analysis agents leverage domain knowledge to gather diverse information from the environment, subsequently storing it in the memory module. The memory module utilizes a memory allocator to distribute the acquired information across memory layers of varying depths. Subsequently, the most pertinent information for the current decision is placed into working memory, where dual decision agents output trading actions. Finally, multi-timescale rewards guide reflection, storing experiential knowledge into deeper memory layers.
4.1. Market Signal Processing and Analysis
To ensure task clarity and enable LLMs to perform effectively in complex market environments, FinPos draws inspiration from institutional investment workflows. It assigns specialized agents to distinct information domains—such as macroeconomics, industry trends, and corporate fundamentals—forming a division of labor similar to that in hedge funds or private equity firms. For each domain, we further establish a two-tier structure: Signal Processing Agents and Analysis Agents.
Signal Processing Agents: These agents are responsible for preprocessing and filtering high-noise, heterogeneous market data streams. We observe that sentiment analysis algorithms often produce extreme or even misleading judgments; therefore, leveraging LLMs’ superior language understanding for data filtering proves more reliable. Through domain-specific prompts and heuristic rules, the agents clean, compress, and prioritize information by relevance and importance. For example, large volumes of low-value or weakly related news are downweighted or discarded, while impactful macroeconomic policies or firm-level events are highlighted and passed on to subsequent modules.
Analysis Agents: These agents operate on the filtered data. Each is designed to address a specific information domain and is enhanced with relevant financial knowledge. We observe that large language models often exhibit financial hallucinations in market scenarios, making it difficult for them to accurately capture the key factors influencing stock prices. They tend to rely solely on explicit keywords in news while overlooking the underlying market logic. For example, an LLM may fail to recognize that news about the S&P 500 is strongly correlated with the performance of leading U.S. stocks. To mitigate this issue, we explicitly inject financial knowledge into the prompts, thereby strengthening their causal awareness. As financial reasoning is gradually integrated into the prompts, the agents evolve from mechanical summarizers into analysts with genuine financial reasoning ability, capable of generating deeper insights and producing more accurate interpretations of market dynamics.
Subsequently, as illustrated in Fig. 2, all analytical results are aggregated into a Hierarchical Memory Module. In this module, important long-term information (e.g., annual reports) is allocated to deep memory, while volatile short-term information (e.g., corporate news) is stored in shallow memory. The hierarchy of memory is not static but dynamically adjusted through post-decision reflection: memories that repeatedly prove their validity are gradually migrated into deeper layers, thereby increasing their weight in future decision-making.
4.2. Dual Trading Decision
In real-world trading, an account’s current holdings directly reflect its risk exposure. Professional traders typically adjust their buy–sell decisions dynamically based on existing positions to maintain a relative balance between risk and return. Following this principle, FinPos adopts a dual-agent framework that embeds position awareness into both the Direction Decision Agent and the Quantity and Risk Decision Agent. By integrating multi-source information retrieved from the memory layer, the dual-agent framework enables a decision-making mechanism that more closely mirrors practical trading logic.
4.2.1. The Direction Decision Agent
Incorporating position awareness requires integrating longer-term information. Unlike conventional direction agents that only focus on individual trades, our agent combines structured memory (news, financial reports, and past actions) with the current portfolio state and explicitly specifies the strategic intent of each action—whether it represents a long-term position-building move or a short-term tactical adjustment to exploit local trends. This explicit articulation not only enhances interpretability for human supervisors but also provides downstream quantity agents with a clear strategic context. To prevent the agent from over-relying on short-term fluctuations, we design prompts that encourage it to factor in longer horizons, thereby gradually cultivating its long-term planning capability.
4.2.2. Quantity and Risk Decision Agent
Most existing trading agents assume fixed-unit trades, overlooking the dynamic adjustment of position size and risk exposure. To overcome this limitation, FinPos introduces the Quantity and Risk Decision Agent, which determines the specific trade size after a directional decision has been made. This agent considers the account’s position state, structured memory, and risk references (e.g. Conditional Value at Risk, CVaR Rockafellar et al. (2000); see Appendix B.1) to exercise more precise control over capital allocation and risk levels. In addition, to prevent excessive exposure, we cap the size of each trade to the maximum suggested by the 95% CVaR. Such a mechanism improves adaptability under market volatility and supports balancing short-term opportunities with long-term robustness.
4.3. Multi-Timescale Reward Design
To guide the agent toward non-myopic trading behavior, we design a trend-aware reward function that leverages multi-timescale market signals. Specifically, we calculate three future price trends in each timestep $t$. $M_{t}^{s}=price[t+1]-price[t]$: 1-day trend (short-term), $M_{t}^{m}=price[t+7]-price[t]$: 7-day trend (mid-term), $M_{t}^{l}=price[t+30]-price[t]$: 30-day trend (long-term). We define the multi-timescale score $M_{t}$ as:
$M_{t}=M_{t}^{s}+M_{t}^{m}+M_{t}^{l}$ (3)
This score represents the aggregated expected trend across multiple timescale (e.g., 1-day, 7-day, and 30-day). The reward at time $t$, $Reward_{t}$ is defined as follows:
[…]
$position_{t}=position_{t-1}+d_{t}\times q_{t}$ (5)
$d_{t}$ and $q_{t}$ denote the direction and quantity of the purchasing decision at the current time step, respectively.
This reward design serves two purposes: 1. Encourage trend alignment: The agent receives positive reinforcement when its predicted direction matches the actual market movement. In this way, the agent is guided away from overfitting to daily fluctuations and instead learns to associate price trajectories with textual information and broader market trends, enabling more robust causal inference in decision-making. 2. Penalty for passivity: In our experiments, we find that during periods of high volatility, agents tend to maintain a inactive state. To address this, we introduce a quadratic penalty on $|M_{t}|$. whenever the position remains unchanged. This discourages missed opportunities in volatile markets while also preventing excessive trading in stable periods.
5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
Our research utilizes actual financial data sourced from publicly authoritative providers and encompasses various forms of market information. (1) Stock prices, obtained from Yahoo Finance, include daily open-high-low-close-volume (OHLCV) data. (2) Company news, retrieved using the Finnhub stock API, includes each article’s related company name, publication date, headline, and summary text, processed for sentiment and semantic relevance. (3) Macroeconomic news, also retrieved via the Finnhub stock API, contains publication dates, headlines, and summary texts for each article. (4) 10-Q (quarterly reports) and 10-K (annual reports), accessed through the U.S. Securities and Exchange Commission (SEC) EDGAR API. These documents provide financial information that publicly listed companies are required to disclose, and standardized into a daily time series format.
5.1.2. Comparative Methods
To assess the effectiveness of FinPos, we conduct comparative experiments with four LLM-based trading agents: FinGPT Yang et al. (2023), FINMEM Yu et al. (2024a), FinAgent Zhang et al. (2024b), and FINCON Yu et al. (2024b); three deep reinforcement learning (DRL) methods (A2C, PPO, DQN); two rule-based methods (MACD, RSI); and two market baselines (Random, Buy-and-Hold). All compared models are evaluated using identical data splits and executed under the same market conditions.
5.1.3. Evaluation Metrics
We evaluate performance using cumulative return (CR%), sharpe ratio (SR), and maximum drawdown (MDD%). CR reflects overall profitability, while SR measures the risk-adjusted return by showing how much excess return it generates per unit of total risk. MDD quantifies the worst-case loss over the trading horizon, which in the Position-Aware Trading Task directly reflects exposure and vulnerability under held positions. Since CR has been introduced earlier (see Eq. (2)), here we provide the formulas for SR and MDD as follows:
$\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}},\quad$ (6)
where $R_{p}$ is the average return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma_{p}$ is the standard deviation of portfolio returns. A higher sharpe ratio indicates more efficient risk-adjusted performance.
$\text{MDD}=\max_{t=1}^{N}\left(\frac{P^{t}-P_{trough}^{t}}{P^{t}}\right)$ (7)
where $t$ denotes the index of the trading day, $P^{t}$ is the account value (the market value of the current stock position) at day $t$, and $P_{trough}^{t}$ is the lowest future account value observed after day $t$. A smaller MDD reflects stronger downside protection and greater robustness of the strategy.
5.1.4. Implementation Details
All LLM trading agents are deployed using GPT-4o Hurst et al. (2024) with a temperature setting of 0.7. All models are evaluated under the Position-Aware Trading Task formally defined in Sec. 3.2, ensuring profits and risks are calculated with explicit holding dynamics. Training covered Jan 2024–Feb 2025; testing spanned Mar–Apr 2025, coinciding with the U.S. election and other major macro events, which makes it a representative and challenging evaluation period.
Table 1. Performance comparison of different models on five stocks
  Models TSLA AAPL AMZN NFLX COIN
CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓
[…]
  Random 32.13-32.13 0.91-0.91 $62.10$ 34.21-34.21 0.28-0.28 $59.40$ 52.35-52.35 0.32-0.32 $61.20$ 1.04-1.04 0.02-0.02 $22.60$ 0.22-0.22 0.01-0.01 $17.60$
Buy and Hold Feldman et al. (2015) 16.40-16.40 0.71-0.71 $63.00$ 16.03-16.03 0.54-0.54 $47.94$ 15.53-15.53 0.45-0.45 $68.91$ $6.84$ $0.42$ $23.08$ 7.65-7.65 0.01-0.01 $16.73$
[…]
  MACD Wang and Kim (2018) 49.21-49.21 0.63-0.63 $74.20$ 46.83-46.83 0.57-0.57 $61.40$ 44.92-44.92 0.49-0.49 $70.30$ 7.42-7.42 0.18-0.18 $27.90$ 9.87-9.87 0.21-0.21 $19.80$
RSI Belafsky et al. (2002) 46.13-46.13 0.59-0.59 $72.80$ 44.28-44.28 0.54-0.54 $59.70$ 42.11-42.11 0.46-0.46 $68.20$ 6.98-6.98 0.16-0.16 $26.70$ 8.65-8.65 0.19-0.19 $19.30$
[…]
  A3C Kang et al. (2018) 92.71-92.71 1.28-1.28 $96.60$ 81.55-81.55 1.14-1.14 $73.40$ 58.24-58.24 0.61-0.61 $72.80$ 5.24-5.24 0.12-0.12 $23.60$ 6.10-6.10 0.09-0.09 $18.60$
DQN Jeong and Kim (2019) 78.25-78.25 0.77-0.77 $80.62$ 72.48-72.48 0.92-0.92 $70.30$ 52.77-52.77 0.48-0.48 $68.10$ $0.78$ $0.08$ $22.10$ 4.25-4.25 0.06-0.06 $18.20$
PPO Li et al. (2025) 73.76-73.76 0.56-0.56 $69.88$ 58.33-58.33 0.55-0.55 $61.40$ 43.12-43.12 0.39-0.39 $64.90$ $1.72$ $0.02$ $21.30$ 2.04-2.04 0.04-0.04 $17.80$
[…]
  FinGPT Yang et al. (2023) 95.19-95.19 1.07-1.07 $94.36$ 85.22-85.22 1.52-1.52 $74.60$ 56.87-56.87 0.54-0.54 $70.30$ $2.01$ $0.01$ $22.30$ 3.63-3.63 0.04-0.04 $17.00$
FinAgent Zhang et al. (2024b) 74.31-74.31 0.89-0.89 $85.65$ 78.00-78.00 1.09-1.09 $71.50$ 49.64-49.64 0.53-0.53 $61.20$ $1.12$ $0.02$ $21.50$ 2.10-2.10 0.03-0.03 $17.50$
FINMEM Yu et al. (2024a) 44.03-44.03 0.52-0.52 $72.10$ 45.88-45.88 0.54-0.54 $47.80$ 41.21-41.21 0.36-0.36 $66.10$ $0.41$ $0.05$ $25.20$ $0.13$ $0.01$ $19.20$
FINCON Yu et al. (2024b) $7.76$ $0.38$ $59.13$ 16.02-16.02 0.13-0.13 $29.03$ 3.30-3.30 0.02-0.02 $54.77$ $7.92$ $0.63$ $20.70$ $7.35$ 1.03 $15.10$
[…]
5.2. Main Results
We evaluated FinPos on five representative stocks—Tesla Inc. (TSLA, electric vehicles), Apple Inc. (AAPL, consumer technology), Amazon Inc. (AMZN, e-commerce), Netflix Inc. (NFLX, online entertainment), and Coinbase Global, Inc. (COIN, cryptocurrency trading). These assets span diverse sectors and exhibit distinct volatility characteristics during the test period: TSLA experienced the highest turbulence, AAPL and AMZN showed relatively stable dynamics, while NFLX and COIN displayed only mild fluctuations.
As shown in Tab 1, FinPos consistently outperforms all baselines, achieving higher CR, superior SR, and lower MDD across all five stocks. For high-volatility assets such as TSLA, AAPL, and AMZN, FinPos demonstrates remarkable robustness. While traditional DRL baselines such as A3C, DQN, and PPO all generated negative returns (e.g., A3C lost over 90% on TSLA with an MDD of 96.6%), and other LLM agents like FinGPT and FinAgent also suffered severe drawdowns (both exceeding 85% on TSLA), FinPos consistently maintained positive performance. Specifically, FinPos achieved CRs of 54.99%, 60.28%, and 52.10% on TSLA, AAPL, and AMZN respectively, while keeping MDDs controlled at 42.34%, 18.44%, and 36.92%. These results show that FinPos not only avoided catastrophic losses faced by DRL and LLM baselines, but also delivered sustained growth exceeding 50% on average under turbulent market conditions. For relatively stable assets such as NFLX and COIN, FinPos also delivered clear advantages. On NFLX, while DRL baselines hovered around neutral or slightly negative returns (e.g., PPO: 1.72%, DQN: 0.78%, A3C: -5.24%), FinPos achieved a CR of 16.85% with an SR of 0.87 and an MDD limited to 20.05%. Similarly, on COIN, FinPos obtained a CR of 14.74% , more than doubling the performance of LLM baselines such as FinCON (7.35%), while keeping downside risk under control (MDD 15.10%). These results demonstrate that FinPos maintains efficiency and steady growth even in calmer markets, combining superior profitability with stable risk management. This directly demonstrates the effectiveness of FinPos’s position-aware and risk-sensitive design.
5.3. Ablation Studies
We evaluated the contribution of each FinPos component through ablation experiments on three representative assets (TSLA, AAPL, AMZN).
  • Position Awareness (PA): Injects explicit position awareness into the agent by equipping the trading agent with a Quantity Decision module. Removing it turns the model into a position-agnostic agent, unable to condition its actions on current holdings.
  • Financial Insight Prompting (FIP): A targeted prompting strategy designed to mitigate LLMs’ weaknesses in financial reasoning. It gradually instills financial thinking by emphasizing causal chains, market trend, and probabilistic inference.
  • Market Signal Processing (MSP): This module consists of multiple agents, each responsible for structuring and filtering raw data from different sources—including financial reports, news, and macroeconomic indicators—to provide coherent and refined market insights.
  • Multi-Timescale Reward (MTR): A self-reflection mechanism that enforces decision consistency across timescales without relying on external reward models.
We select TSLA, AAPL, and AMZN as test assets due to their rich set of news and events. Tab. 2 reports results under different module removals, showing that the full model consistently achieves the best CR, SR, and MDD. This demonstrates that FinPos’s various modules operate synergistically to yield optimal trading decisions.
Table 2. Ablation Results on TSLA, AAPL and AMZN
PA FIP MSP MTR TSLA AAPL AMZN
CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓ CR%↑ SR↑ MDD%↓
39.19 0.45 82.89 45.96 0.49 36.10 36.98 0.43 73.71
41.57 0.49 48.78 48.57 0.54 21.24 39.38 0.50 41.76
44.35 0.55 53.27 50.64 0.59 24.25 42.19 0.56 47.57
48.68 0.59 59.07 57.36 0.66 25.73 47.24 0.60 51.50
54.99 0.67 42.34 60.28 0.69 18.44 52.10 0.64 36.92
5.3.1. Position Awareness (PA)
As introduced in Sec. 4.2, the Trading Decision Module equips FinPos with position awareness. The ablation results in Tab. 2 confirm its necessity: on TSLA, CR drops from 54.99% to 39.19%, SR falls from 0.67 to 0.45, and MDD nearly doubles from 42.34% to 82.89% once PA is removed. Similar degradations appear on AAPL and AMZN. These findings validate that PA is indispensable for translating direction forecasts into position-based execution, ensuring both profitability and stability.
Figure 3. We conducted a fine-grained analysis of agent prompts along the dimensions of eight characteristics and three levels of emphasis.
Figure 4. The impact of varying the maximum timescale of the multi-timescale reward on various performance metrics.
5.3.2. Financial Insight Prompting (FIP)
As shown in Tab. 2, introducing FIP leads to a clear overall performance gain, consistently enhancing cumulative return across all assets (e.g., TSLA: 41.57% → 54.99%; AAPL: 48.57% → 60.28%). In our ablation hierarchy, FIP ranks second only to the position-aware mechanism in terms of its contribution to overall performance.
To investigate the impact of prompt design on the depth of agent finance insight, we divide the content of prompts into eight key dimensions: Role Setting, Task Breakdown, Depth Analysis, Multi-Dimensional Analysis, Logical Reasoning, Contextual Consideration, Actionability, and Clarity & Conciseness(the detailed explanations of these eight dimensions are provided in Appendix C.2.1). For each dimension, we design prompts with different levels of refinement—Low, Medium, and High—and conduct graded experiments. The results are shown in Fig. 4.
The results indicate that Task Breakdown, Depth Analysis, and Multi-dimensional Analysis are the most influential dimensions for overall agent performance. Notably, Depth Analysis plays a particularly crucial role in financial reasoning: since the quality of financial analysis is significantly influenced by deep insights, particularly those concerning causal chains and risk-reward tradeoffs. In contrast, Multi-Dimensional Analysis presents a more nuanced pattern. Financial analysis naturally spans across multiple perspectives, such as macro vs. micro and short-term vs. long-term. While moderate prompting enhances the model’s ability to cover multiple factors, overly elaborate prompts may overwhelm the LLM, causing it to “overthink,” lose focus, and eventually degrade performance—an effect we refer to as the information burden. Among the eight dimensions, Contextual Consideration shows the most pronounced performance shift. Without explicit prompting, the model tends to overlook this dimension—for instance, it may treat tariff adjustments as entirely unrelated to stock prices. Once guided, however, the agent develops stronger contextual reasoning: it can link major events (e.g., political shifts, regulatory changes) to firm-level responses (e.g., strategic adjustments to tariffs), and further to market outcomes. This ability allows the model to generate insights that align more closely with real-world market dynamics.
Overall, the optimal strategy in FIP design is to selectively emphasize key dimensions—particularly Depth Analysis, Contextual Consideration, and Logical Reasoning—while keeping the prompt concise and focused. Such design not only improves the quality of reasoning but also ensures that the conclusions provide actionable guidance in real financial markets.
5.3.3. Market Signal Processing (MSP) Module
We compare FinPos with a variant that directly feeds unfiltered raw data into the analysis agents. Removing MSP results in a consistent decline across all performance metrics (see Tab. 2). For instance, on TSLA, the cumulative return drops from 54.99% to 44.35%, the Sharpe Ratio drops from 0.67 to 0.55, and the maximum drawdown increases from 42.34% to 53.27%. Similar degradations are observed on AAPL (CR: 60.28% → 50.64%; SR: 0.69 → 0.59; MDD: 18.44% → 24.25%) and AMZN (CR: 52.10% → 42.19%; SR: 0.64 → 0.56; MDD: 36.92% → 47.57%). These results indicate that without structured filtering, the downstream analysis agents struggle to identify key signals from noisy textual inputs, leading to delayed or inaccurate trading responses. By contrast, the full FinPos system—with MSP-enabled preprocessing—produces cleaner and more coherent insights, thereby enhancing both trading precision and overall robustness.
5.3.4. Multi-Timescale Reward (MTR)
Finpos self-reflects its decision based on returns across multiple time scales (Sec. 4.3). Our experiments show that MTR is crucial for the agent’s long-term risk control. Without it, the agent loses sensitivity to long-horizon risks, leading to a sharp deterioration in drawdown management. As shown in Tab. 2, on TSLA, CR drops from 54.99% to 48.68%, SR decreases from 0.67 to 0.59, and MDD surges from 42.34% to 59.07%—a clear sign that the lack of multi-timescale self-reflection severely weakens the agent’s ability to capture stable trends and risk control. Similar degradations are observed on AAPL (CR: 60.28% → 57.36%; MDD: 18.44% → 25.73%) and AMZN (CR: 52.10% → 47.24%; MDD: 36.92% → 51.50%).
The effectiveness of MTR depends on the timescale used to compute the overall reward. In our design, the MTR module incorporates three reward horizons (1, 7, and 30 days) to jointly capture immediate fluctuations, medium-term trends, and long-term trends. Among these, the long-term trends serves as the principal timescale that governs the agent’s reflection and strategic consistency. Therefore, our sensitivity study focuses on the long-term trends, varying its length to examine the trade-off between adaptiveness and stability. As shown in Fig. 4, very short horizons (7–14 days) yield weak performance, as the agent focuses excessively on local fluctuations without forming stable strategies. Extending the horizon improves consistency, with 30 days emerging as the optimal setting: CR and SR peak, while drawdowns are minimized. However, when the horizon exceeds 30 days, performance declines again—longer windows dilute the reward signal and hinder the reflection process, making the agent less adaptive to new market conditions. This sensitivity study highlights that MTR is most effective when balancing short-term responsiveness with long-term stability, and in our experiments, 30 days provides the best trade-off.
5.4. Risk Analysis
Beyond numerical results, Fig. 5 provides complementary evidence from profitability and risk control. The top-left panel reports the Calmar ratio (see Appendix B.2), which directly reflects risk-adjusted performance under extreme losses. A higher value thus indicates that the strategy achieves strong profitability while effectively controlling downside risk. Here, FinPos achieves the highest score, suggesting that its profitability is supported by effective drawdown management rather than excessive risk-taking. The bottom-left panel plots cumulative return (CR %) against risk-control ability (1 - MDD). FinPos is positioned at the upper-right frontier, combining high profitability with robust risk management, whereas most baselines cluster in regions of either low return or weak control. On the right, the time-series curves and the exposure-risk overlay further illustrate the critical role of position-aware management: around high-volatility events (vertical dashed lines), methods without PA suffer sharp exposure spikes and subsequent drawdowns, while FinPos is capable of detecting the onset of risks and implementing defensive maneuvers. Taken together, these three views consistently demonstrate that position awareness not only mitigates short-term shocks but also improves risk-adjusted outcomes over the entire horizon. Additional analyses for other stocks are provided in Appendix C.1
Figure 5. Profitability versus risk dynamics on TSLA. Top-left: Calmar ratio, capturing return relative to maximum drawdown (higher is better). Bottom-left: joint view of profitability (CR%) and risk-control (1 - MDD), indicating return–risk balance. Right: time-series comparison of cumulative return (bottom) and exposure risk (top) across major events. Vertical dashed lines mark the occurrence of major events.
6. Conclusion
In this paper, we introduce a position-aware trading task that more closely resembles real market conditions compared to the single-step trading task previously employed by researchers. To address this task, we design FinPos, a novel LLM Trading Agent equipped with position-awareness and risk management capabilities. To meet the demands of position management, an LLM agent must possess heightened market sensitivity, enhanced risk control abilities, and a longer-term perspective. Consequently, FinPos employs a specialized market perception module, a dual-decision agent system, and a multi-timescale reward design to achieve these objectives. Our experiments demonstrate that the position management task imposes higher comprehensive ability requirements on LLM Agents. FinPos explores solutions for trading agents operating in this more realistic market environment. However, we also identify two limitations of FinPos that require further attention. Firstly, FinPos is unable to implement coordinated management of multiple asset allocations, thus lacking the capability for higher-level operations such as risk hedging. Secondly, FinPos has been tested only on stock-type investment assets, and its applicability to other asset classes such as futures and bonds remains to be validated.
Appendix A The implementation details of FinPos
To provide a clearer understanding of how FinPos operates in practice, we include the complete prompt designs used by each module. These prompts define how the system processes financial information, reasons about market dynamics, and executes trading decisions in a structured manner.
A.1. Market Signal Processing and Analysis Module
This module provides the detailed prompts used in the Market Signal Processing and Analysis Module. Each category of textual data (financial reports, quarterly reports, company news, and macroeconomic news) is associated with two prompt templates: one for extracting sentiment and insight, and the other for assessing risk and potential market impact.
A.1.1. 10-K Filings (Annual Reports)
This section presents the prompts used to process and analyze annual 10-K reports. The first prompt extracts and filters key financial information, while the second evaluates its implications for the company’s stock price.
System Prompt: You are a professional financial analyst. Your task is to evaluate the potential impact of the key points extracted from a companys 10-K report on its stock price, providing decision support for an intelligent trading agent.
You have received the following 10-K key points about {symbol}:
"{filtered_key_points}"
Target company: {symbol}
Please complete the following tasks:
1. Screen and extract the most critical financial and operational highlights, focusing on major changes, performance deviations, strategic shifts, and newly emerging risks or opportunities.
2. Filter out repetitive, boilerplate, or investor-irrelevant information.
3. Rank the retained key points by their importance to investors.
response_format_prompt = Please return the result in the following JSON format, without adding any other explanation:
"key_points": "Selected key points, sorted by importance."
"reason": "An explanation of why these key points were retained and the rationale behind their importance over other content."
System Prompt: You are a professional financial analyst specializing in evaluating the short-term and medium-to-long-term impacts of company 10-K reports on stock prices. You are providing decision support for an intelligent trading agent.
You have received a summary of the annual 10-K report for {symbol}:
"{agent_scratch}"
Target company: {symbol}
1. Analyze the potential **short-term (days to one week)** and **medium-to-long-term (weeks to months)** effects on {symbol}’s stock price. Consider whether the developments are likely to surprise the market **positively or negatively** based on typical investor expectations and sentiment.
2. For each impact direction, **differentiate** between contributing factors (e.g., profitability, cash flow, capital allocation, competitive positioning, regulatory risk). Analyze **interactions or trade-offs** between opposing forces.
3. Explain your reasoning in a structured, multi-dimensional way. Go beyond summarization-**synthesize the data**, explore **counterfactual scenarios**, and account for **macro and industry context**. If relevant, mention **investor psychology** or narrative shifts.
4. DO NOT simply restate the report. Your goal is to **interpret, evaluate, and draw meaningful implications** for trading behavior and valuation outlook.
Please maintain professionalism, clarity, and logical coherence. Highlight key opportunities and risks with balanced and nuanced judgment.
Please output only the response in JSON format without any additional commentary:
{response_format_prompt}= Please return the result in the following JSON format, without adding any other explanation:
"insight": "This 10-K report is positive/negative/neutral for {symbol} in the short term, and positive/negative/neutral in the medium to long term.",
"reason": "Explain the core reasoning behind the judgment, reflecting logical analysis of the key points."
A.1.2. 10-Q Filings (Quarterly Reports)
Similarly, the following prompts are designed for 10-Q reports, focusing on quarterly performance updates and market reactions.
System Prompt: You are a professional financial analyst. Your task is to evaluate the potential impact of the key points extracted from a companys 10-Q report on its stock price, providing decision support for an intelligent trading agent.
You have received the following 10-Q key points about {symbol}:
"{filtered_key_points}"
Target company: {symbol}
Please complete the following tasks:
1. Screen and extract the most critical financial and operational highlights, focusing on major changes, performance deviations, strategic shifts, and newly emerging risks or opportunities.
2. Filter out repetitive, boilerplate, or investor-irrelevant information.
3. Rank the retained key points by their importance to investors.
response_format_prompt = Please return the result in the following JSON format, without adding any other explanation:
"key_points": "Selected key points, sorted by importance."
"reason": "An explanation of why these key points were retained and the rationale behind their importance over other content."
System Prompt: You are a professional financial analyst specializing in evaluating the short-term and medium-to-long-term impacts of company 10-Q reports on stock prices. You are providing decision support for an intelligent trading agent.
You have received a summary of the quarterly 10-Q report for {symbol}:
"{agent_scratch}"
Target company: {symbol}
1. Analyze the potential **short-term (days to one week)** and **medium-to-long-term (weeks to months)** effects on {symbol}’s stock price. Consider whether the developments are likely to surprise the market **positively or negatively** based on typical investor expectations and sentiment.
2. For each impact direction, **differentiate** between contributing factors (e.g., profitability, cash flow, capital allocation, competitive positioning, regulatory risk). Analyze **interactions or trade-offs** between opposing forces.
3. Explain your reasoning in a structured, multi-dimensional way. Go beyond summarization-**synthesize the data**, explore **counterfactual scenarios**, and account for **macro and industry context**. If relevant, mention **investor psychology** or narrative shifts.
4. DO NOT simply restate the report. Your goal is to **interpret, evaluate, and draw meaningful implications** for trading behavior and valuation outlook.
Please maintain professionalism, clarity, and logical coherence. Highlight key opportunities and risks with balanced and nuanced judgment.
Please output only the response in JSON format without any additional commentary:
{response_format_prompt}
response_format_prompt = """Please respond in the following JSON format **without adding any additional explanations**:
"key_points": "Concise summary of the most critical content, including key highlights and risks with brief explanation.",
"insight": "This report has a positive/negative/neutral impact on {symbol} in the short term, and a positive/negative/neutral impact in the medium to long term.",
"reason": "Comprehensive explanation of the reasoning behind the judgment, showing multi-dimensional logical analysis and complex factor consideration rather than simple summary."
A.1.3. Macroeconomic News
This section presents the prompts used to process macroeconomic news and policy releases. Two corresponding prompt templates are adopted: the first filters and ranks macroeconomic news items by their relevance and significance to thethe target company, while the second analyzes how these events may influence investor sentiment, capital flows, and asset price movements over short and medium-to-long horizons.
System Prompt: You are an experienced financial research assistant. Your task is to determine whether a given news article is related to a specific company.
Target company: {symbol}
Please analyze the news below and classify the relationship between the news and the company as either "direct", "indirect", or "none", according to the criteria provided:
Classification criteria:
1. If the news **explicitly mentions** the company name (e.g., Tesla), its executives (e.g., Elon Musk), its products, financial reports, mergers, partnerships, or investments -- classify as **direct**
2. If the news does **not explicitly mention** the company, but includes topics that **have a substantial impact on the companys business, valuation, or market performance** -- classify as **indirect**, such as:
- Industry level: industry trends, changes in market demand, technological advancements, industry regulatory policies, upstream/downstream supply chain, competitor dynamics, price fluctuations of key materials (e.g., lithium, batteries)
- Macroeconomic factors: macroeconomy, Federal Reserve policies, interest rates, inflation, employment, consumer spending, GDP growth, manufacturing indices, PMI, retail sales, and other macroeconomic indicators
- Financial market sentiment: significant fluctuations or sustained trends in the S&P 500, Nasdaq, Dow Jones indices; market overheating, overbought conditions, or panic selling that may affect overall risk appetite; valuation adjustments in tech/growth stock sectors; market rotation; financing environment; IPO activities; large ETF inflows or outflows
- Policies and regulations: national policies, taxation, regulation, energy, climate, green transition, green energy subsidies, emission standards, electric vehicle regulations; US-China trade war, export restrictions, chip bans, customs policies, etc.
- US-China relations, export controls, trade wars, tariff adjustments, technology bans
- Geopolitical conflicts (e.g., Russia-Ukraine war, Middle East tensions), international sanctions, energy price surges causing global market volatility or disruptions in energy/logistics/supply chains
- Key figures (e.g., Trump, Biden, Federal Reserve Chair Powell) making political, economic, or policy statements, policy preferences, election outlooks, trade comments, or antitrust remarks
3. If the news is **not substantively related** to the company and is **unlikely to impact** its operations or stock price -- classify as **none**, such as:
- Natural disasters, entertainment gossip, or local events unrelated to the companys business, industry, or market
- Regional incidents with no significant impact on the companys countrys economy or policies
News article:
{agent_scratch}
Please output only the response in JSON format without any additional commentary:
{response_format_prompt}
response_format_prompt = """Please output a JSON in the following format:
"relation_type": "direct" | "indirect" | "none",
"reason": "Briefly explain the reasoning behind your judgment"
System Prompt: You are a professional financial analyst specializing in evaluating the medium- to long-term impact of financial news on company stock prices. You are assisting an intelligent trading agent with decision-making support.
You have received the following financial news:
"{agent_scratch}"
The target company is: {symbol}
Please complete the following tasks:
1. Do **not** repeat or summarize the original news content;
2. Determine whether this news has a **material impact** on {symbol}’s stock price, not limited to direct relevance - please also consider macroeconomic policy, supply chain dynamics, market sentiment, geopolitical risks, or other indirect or lagging factors;
3. If there is an impact, provide **one clear and concise investment insight**, explaining how the news might affect {symbol}’s stock price in the coming **weeks to months** (e.g., bullish or bearish);
4. If there is **no clear relevance or impact**, clearly state that the news has **no significant effect** on {symbol};
5. Evaluate the relevance level of the news to {symbol}, using the following scale:
"high": The news has a direct and significant impact on the companys fundamentals, financials, regulatory environment, or industry position;
"medium": The news could have an indirect or delayed impact, such as through macroeconomic trends, industry supply/demand shifts, investor sentiment, or cost structure changes;
"low": The news is largely unrelated or only remotely connected to the company.
Please output only the response in JSON format without any additional commentary:
{response_format_prompt}
response_format_prompt = """Please respond using the following JSON format and do not include any additional text:
"insight": "Summary of how this news may impact {symbol}",
"relevance": "high" | "medium" | "low"
A.1.4. company news
Since company-specific news can be directly collected by ticker symbol, the filtering process focuses primarily on assessing **relevance, materiality, and potential market impact**, rather than broad topic association. Two prompt templates are used in this module - one for filtering and ranking important news items, and the other for analyzing their short-term and medium-to-long-term effects on stock performance.
System Prompt: You are a professional financial analyst. Your task is to filter and prioritize firm-level news items based on their potential importance to investors and their relevance to the companys stock price.
You have received several pieces of company-related news for {symbol}:
"{news_batch}"
Please complete the following steps:
1. Identify which items are **material** and likely to influence investor perception or price movement;
2. Filter out minor, repetitive, or purely descriptive updates with limited market relevance;
3. Rank the retained items by their expected significance to the stock price, considering tone, topic, and potential investor reaction.
Please return the result strictly in JSON format:
"key_points": "Selected and ranked company news items that are most likely to affect {symbol}’s stock price.",
"reason": "Explain briefly why these items are more significant than others."
System Prompt: You are a professional financial analyst specializing in assessing the **price sensitivity** of company-related news. You are assisting a high-performance trading agent that only acts based on material, relevant information.
Here is a piece of news you received:
"{agent_scratch}"
Target company: {symbol}
Please follow these instructions:
1. Do NOT summarize the news content;
2. Focus ONLY on the potential impact of this news on {symbol}’s stock price;
3. If this news is irrelevant or has no clear directional impact on {symbol}, clearly mark it as **"neutral"** with an appropriate reason;
4. Evaluate the likely impact in both:
- **Short term** (1-5 trading days)
- **Medium to long term** (a few weeks to months);
5. Be strict: only assign "positive" or "negative" if the news provides clear evidence of directional influence on {symbol}’s fundamentals or investor sentiment.
Please output only the response in JSON format without any additional commentary:
response_format_prompt = """Please return the result in the following **JSON format**, without adding any extra explanation:
"insight": "This news has a [positive/negative/neutral] impact on {symbol} in the short term, and a [positive/negative/neutral] impact in the medium to long term.",
"reason": "Explain the key reasoning behind your assessment. Do not summarize the news content."
A.2. Dual Trading Decision Module
This section provides the detailed prompt structures used in the Dual Trading Decision Module. While the architectural overview (see Fig. 2) already explains the interaction flow, here we focus on the internal prompt logic and reasoning objectives of each decision agent.
A.2.1. Direction Decision Agent
The following prompt guides the agent to leverage the key insights extracted by the preceding analytical modules and determine the optimal trading direction (buy, sell, or hold), along with the overall strategic orientation for the current trade.
System Prompt:
# memory IDs
short_memory_id_desc = "ID of short-term information."
mid_memory_id_desc = "ID of mid-term information."
long_memory_id_desc = "ID of long-term information."
reflection_memory_id_desc = "ID of reflective-period information."
train_memory_id_extract_prompt = "Select and store the most investment-relevant information from major sources (e.g., ARK, Two Sigma, Bridgewater Associates) into the {memory_layer} memory."
test_memory_id_extract_prompt = "Retrieve the most relevant information from the {memory_layer} memory for the current investment decision."
# trading summary
train_trade_reason_summary = "Based on a professional traders advice, explain why the trader would make such a decision given the provided information."
test_trade_reason_summary = "Based on the text information and summarized price trends, explain the reason for your investment decision."
test_invest_action_choice = "Based on the information, choose one of the following actions: buy, sell, or hold."
# investment info
train_investment_info_prefix = (
"The current date is {cur_date}. The observed market facts are as follows: "
"For {symbol}, the price difference between the next and current trading day is {cur_record_t1}; "
"the 7-day difference is {cur_record_t7}; "
"the 30-day difference is {cur_record_t30}. "
[…]
test_investment_info_prefix = "The stock under analysis is {symbol}, and the current date is {cur_date}."
# sentiment & momentum explanation
test_sentiment_explanation = """For example, positive news about a company may boost investor confidence and trigger buying activities, pushing the stock price upward;
whereas negative news tends to dampen sentiment, leading to selling pressure and price declines.
Additionally, news related to competitors or the broader industry can indirectly affect the target stocks performance.
Sentiment scores (positive, neutral, negative) represent the distribution across these categories (summing to 1) and, together with "importance" and "timeliness" indicators, help assess the market impact and validity of the information.
test_momentum_explanation = """The following summarizes recent price movements, i.e., momentum.
Momentum reflects the idea that stocks performing strongly in the short term often continue rising,
while weak performers are more likely to keep declining.
# training phase prompt
train_prompt = """Please complete the following two tasks based on the investment information below:
Important: Do NOT use any future price differences (T+1, T+7, T+30) in your reasoning. These are unavailable in real-time trading. Any output referencing them will be considered invalid.
1. Directional Decision:
Choose one of the following actions: "buy", "sell", or "hold" (only if uncertain).
You must consider:
- Information from short-, mid-, long-term, and reflective memories;
- Historical price momentum;
- Sentiment tendencies, importance, and timeliness in news or reports.
Briefly describe your decision logic, the overall trading strategy (e.g., long-term accumulation or short-term profit), and indicate the supporting memory indices.
2. Reflection:
The system will automatically evaluate whether your directional judgment matches the market trend.
- If incorrect, explain the misinterpreted or overemphasized information.
- If correct, summarize the key factors behind the correct judgment.
\${investment_info}
Your output must strictly follow the JSON format below, with no extra text:
"investment_decision": "buy" | "sell" | "hold",
"summary_reason": "Brief explanation of your decision logic",
"short_memory_index": [integer list],
"middle_memory_index": [integer list],
"long_memory_index": [integer list],
"reflection_memory_index": [integer list],
"reflection_analysis": "Reflection analysis text"
# testing phase prompt
test_prompt = """Determine the optimal investment direction based on the following information and briefly justify your reasoning.
You must consider:
- Information from all memory layers (short-, mid-, long-term, reflective);
- Historical price momentum;
- The importance, sentiment, and timeliness of key information.
Provide one of three decisions: "buy", "sell", or "hold", and indicate the memory IDs supporting your judgment.
\${investment_info}
\${gr.complete_json_suffix_v2} }
A.2.2. Quantity and Risk Decision Agent
The following prompt instructs the agent to determine the specific order quantity for the current trade based on the analytical results and strategic guidance from the Direction Decision Agent, while adjusting for current holdings and potential risk exposure.
System Prompt:
# memory IDs
short_memory_id_desc = "ID of short-term information."
mid_memory_id_desc = "ID of mid-term information."
long_memory_id_desc = "ID of long-term information."
reflection_memory_id_desc = "ID of reflective-period information."
train_memory_id_extract_prompt = "Select and store the most investment-relevant information from major sources (e.g., ARK, Two Sigma, Bridgewater Associates) into the {memory_layer} memory."
test_memory_id_extract_prompt = "Retrieve the most relevant information from the {memory_layer} memory for the current investment decision."
# trading summary
train_trade_reason_summary = "Based on a professional traders advice, explain why the trader would make such a decision given the provided information."
test_trade_reason_summary = "Based on the text information and summarized price trends, explain the reason for your investment decision."
test_invest_action_choice = "Based on the information, choose one of the following actions: buy, sell, or hold."
# investment info
train_investment_info_prefix = (
"The current date is {cur_date}. The observed market facts are as follows: "
"For {symbol}, the price difference between the next and current trading day is {cur_record_t1}; "
"the 7-day difference is {cur_record_t7}; "
"the 30-day difference is {cur_record_t30}. "
[…]
train_reward_explanation = """Reward reflects the quality of your past decision:
- **Positive**: Good decision; higher means better alignment with market.
- **Negative **: - **Negative**: A weaker decision. The more negative the value, the worse the outcome - may caused by misreading available data.
Use reward **only for reflection**, not for future predictions.
test_investment_info_prefix = "The stock under analysis is {symbol}, and the current date is {cur_date}."
# sentiment & momentum explanation
test_sentiment_explanation = """For example, positive news about a company typically boosts market confidence, stimulates buying, and drives up the stock price;
Negative news, on the other hand, weakens confidence, triggers selling pressure, and causes the stock price to fall.
Industry or competitor dynamics may also indirectly affect the target companys performance.
The sentiment score (positive, neutral, negative) reflects the proportion of the text in each of the three sentiment categories (summing to 1).
It can be combined with the "Importance" and "Timeliness" metrics to assess the market impact and validity of the information.
In addition, you need to combine the output of the previous Direction Decision Agent (i.e., the overall strategic description of this transaction) as the strategic basis for quantitative decisions.
test_momentum_explanation = """The following summarizes recent price movements, i.e., momentum.
Momentum reflects the idea that stocks performing strongly in the short term often continue rising,
while weak performers are more likely to keep declining.
# training phase prompt
train_prompt = """Please complete the following two tasks based on the investment information below:
Important: Do NOT use any future price differences (T+1, T+7, T+30) in your reasoning. These are unavailable in real-time trading. Any output referencing them will be considered invalid.
1. Investment Amount and Risk Decision:
You already know the directional decision (buy/sell/hold) made by the Direction Decision Agent in the previous stage.
Based on this, determine the **specific order quantity**(integer) and ensure that the transaction volume does not exceed the maximum limit {maxcvar} recommended by the risk control module.
You must consider the following factors:
- Information in each memory layer (short-term, medium-term, long-term, and reflection period);
- Historical momentum and price volatility;
- The sentiment, importance, and timeliness of news or financial reports;
- Current account holdings and overall risk exposure;
- Trading strategy determined in the previous phase.
Please briefly explain your quantity decision logic and indicate the memory indexes supporting this decision.
2. Decision Reflection and Analysis:
The system will calculate a reward based on the order quantity and corresponding return.
- If the reward is negative, please explain any market signals or risk factors you may have misjudged;
- If the reward is positive, please summarize the core rationale that led to your correct decision.
\${investment_info}
Your output should strictly adhere to the following JSON format and not include any other content:
"order_size": integer (range 1 to {maxcvar}),
"summary_reason": "Please enter your quantity and risk decision logic here",
"short_memory_index": [list of integers],
"middle_memory_index": [list of integers],
"long_memory_index": [list of integers],
"reflection_memory_index": [list of integers],
"reflection_analysis": "Please fill in your reflection description here."
# Testing phase prompt
test_prompt = """Based on the following information, please determine the **order quantity** for the current trade.
You know the directional decision (buy/sell/hold). Please specify the specific order quantity based on the risk exposure and CVaR constraint (maximum order quantity {maxcvar}).
You must consider:
- Memory information at each level (short-term, medium-term, long-term, reflection period);
- Momentum trend, sentiment, information importance, and timeliness;
- Current account holdings and overall risk;
- Trading strategy for the previous directional decision.
Please output a specific order quantity (integer, not exceeding {maxcvar}) and indicate the information index that supports your judgment.
\${investment_info}
\${gr.complete_json_suffix_v2} }
Appendix B Formulas of Classic Financial Metrics
To better evaluate the risk–return profile of trading strategies, we adopt two widely used risk metrics: the Conditional Value at Risk (CVaR) and the Calmar Ratio. Their formal definitions are as follows.
B.1. Conditional Value at Risk (CVaR)
Let the profit and loss over a trading horizon be denoted by $PnL$. The Value at Risk (VaR) at a confidence level $\alpha$ represents the maximum potential loss not exceeded with probability $\alpha$, formally defined as:
$\text{VaR}_{\alpha}(PnL)=\inf\{l\in\mathbb{R}:P(PnL\leq l)\geq\alpha\}.$
The Conditional Value at Risk (CVaR) measures the expected loss that occurs beyond the VaR threshold, providing a more comprehensive view of downside risk:
$\text{CVaR}_{\alpha}(PnL)=\mathbb{E}[PnL\mid PnL\leq\text{VaR}_{\alpha}(PnL)].$
In our experiments, $PnL$ is calculated based on realized returns over the testing period. A smaller CVaR indicates stronger downside protection and better risk control capability.
B.2. Calmar Ratio
The Calmar Ratio evaluates the trade-off between return and maximum drawdown. It is defined as:
$\text{Calmar Ratio}=\frac{R_{\text{annual}}}{|\text{MDD}|}$
where $R_{\text{annual}}$ denotes the annualized return, and MDD represents the maximum drawdown during the same period. A higher Calmar Ratio indicates a better risk-adjusted performance.
Appendix C More Experiments
Figure 6. Profitability versus risk dynamics. Top-left: Calmar ratio, capturing return relative to maximum drawdown (higher is better). Bottom-left: joint view of profitability (CR%) and risk-control (1 - MDD), indicating return-risk balance. Right: time-series comparison of cumulative return (bottom) and exposure risk (top) across major events; PA-aware strategies mitigate exposure spikes and sustain more stable growth.
C.1. More Stock Trading Result Graphs
Similarly, for AAPL, the supplementary plots (Fig. 6) provide complementary evidence from profitability and risk control. FinPos again occupies the upper-right region in the CR% vs. risk-control space, indicating that its returns are achieved without compromising drawdown management. The time-series and exposure-risk overlays highlight that, during high-volatility periods, non-position-aware methods experience sharp spikes in exposure and subsequent losses, whereas FinPos effectively anticipates and mitigates these risks. Overall, these results reinforce that position awareness consistently improves risk-adjusted performance across different stocks.
C.2. More Details of Ablation Studies
C.2.1. Financial Insight Prompting (FIP)
This section explains the specific contents of the 8 dimensions in the fip section.(Tab. 3)
Table 3. Evaluation Criteria for Prompt Traits
Trait Evaluation Criteria (How to measure)
Role Setting Does it clearly assign a professional role to the AI (e.g., ”Financial Analyst”)?
Task Breakdown Does it break down complex analytical tasks into clear sub-tasks?
Depth Requirement Does it explicitly ask the model to infer and analyze, not just summarize or restate?
Multi-dimensional Analysis Does it guide the model to consider multiple factors (e.g., short-term/long-term, macro/micro)?
Logical Reasoning Does it encourage the model to build causal chains, weigh pros and cons, and infer potential impacts?
Contextual Consideration Does it guide the model to consider non-financial factors (e.g., investor psychology, market narrative)?
Actionability Does it require the model to provide conclusions that have practical guidance for decision-making?
Clarity & Conciseness Is the prompt itself easy to understand, unambiguous, and not redundant?
C.2.2. Macroeconomic News (MN) module
Beyond the experiments mentioned in the main text, FinPOS further broadens its information scope by incorporating macroeconomic news (MN) and other sources into the Market Signal Processing and Analysis module, and conducts corresponding ablation studies. The following provides examples of the decision-making process of the directional agent before and after integrating the macroeconomic news module.
  • Before incorporating MN: I begin by reviewing latest earnings report and recent company news of TSLA. The quarterly results exceed market expectations, and Musk’s announcement about expanding factory investment conveys a strong growth signal. Market sentiment around the company appears to be positive, and technical indicators show sustained upward momentum. Based on the alignment of these signals, I infer that TSLA’s short-term trend is likely to be bullish. Decision: buy.
  • After incorporating MN: Recent news indicates that Trump has introduced new tariffs, inflation is accelerating, and export conditions for U.S. manufacturers are deteriorating. Although TSLA’s fundamentals remain strong, these macro developments imply higher production costs and potential demand pressure. After integrating these broader signals, I reassess the outlook: systemic risk outweighs company-level optimism. Decision: sell.
This comparison demonstrates that MN provides crucial contextual signals for decision-making, especially during periods of high market volatility or systemic risk, significantly enhancing the agent’s risk awareness and decision accuracy.